if A's income is 20% more than B's income, how much percent is B's income less than A's income
Answer by Theo(13342)

Let A = income that A makes.

Let B = income that B makes.

Percent difference in income that A makes compared to B would be equal to (A-B)/B * 100%

Percent difference in income that B makes compared to A would be equal to (B-A)/A * 100%

Assume A makes 120 dollars and B makes 100 dollars.

(A-B)/B = (120-100)/100 = 20/100 = .2 * 100% = 20%

This is equivalent to 20% more than what A makes.

(B-A)/A = (100-120)/120 = -20/120 = -.167 * 100% = -16.7% (rounded)

This is equivalent to 16.7% less than what B makes.

It's the same difference in income ($20).

The denominator is different.

Percent of B's income, then B's income is the denominator.

Percent of A's income, then A's income is the denominator.